OpenStudy (anonymous):

Find two consecutive intergers whose product is 20

OpenStudy (anonymous):

4,5

OpenStudy (anonymous):

You want to find two numbers that sit right next to each other that when multiplied equal 20. For example if the question was 6 instead of 20 the answer would be 2 and 3 Because they sit right next to each other. And 2 times 3 is 6

OpenStudy (anonymous):

yes but how would you solve it ?

OpenStudy (anonymous):

algebraically

OpenStudy (anonymous):

n(n+1) = 20 n^2 + n - 20 = 0 (n + 5)(n - 4) = 0 n = 4 n + 1 = 5 so 4,5

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

20=20*1 20=4*5 20=2*10